OpenStudy (anonymous):

a square pyramid has a base length of 6 inches. the slant height od the pyramid is 8 inches. what is the total surface of the pyramid

OpenStudy (anonymous):

Surface Area equals= A+b=bh+b^2=8(6)+36=84...was that an option on your quiz?

OpenStudy (anonymous):

I got 132?

OpenStudy (anonymous):

i could be wrong..how did you get your answer?